change of time period in a pendulum clock

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
I don’t think your question is complete.
butT = 2 \pi \sqrt{\frac{L}{g}} \,
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty